# Search Timelines: Visualizing Search History to Enable Cross-Session Exploratory Search

### v1

@@ 5.4 RQ2b: Usability During the Resumed Session
The key feature provided to support resuming the search task was the detailed timeline within the workspace. Replicating the search result cards from the SERP into the timeline resulted in a familiar structure in the interface, which contributed to its ease of use. The extra contextual information that linked each saved resource to its source query and temporal information, contributed to perceptions of usefulness and satisfaction. These aspects of usability were further supported by providing the search timeline in an overview format when the searchers returned to issuing queries and viewing search results.
@@ 5.5 RQ3a: Perceived Knowledge Gain During the Initial Search
The differences in the perceived knowledge measured pre-search and post-search during the initial search session were calculated separately for those who used the Baseline and those who used Search Timelines. For both cases, we found an increase in the perceived knowledge that was statistically significant with large effect sizes (see Table 5). This finding shows that the participants of both interfaces were actively engaged in the search process and found search results that felt enhanced their knowledge during the initial search.
We also conducted a statistical analysis comparing the degree of knowledge change between the two interface conditions. Statistical significance with a large effect size was found in favour of Search Timelines ($W=245.5$, p <$\mathbf{10^{-5}}$, $d=2.365$). This outcome highlights the value of providing a visual timeline of queries and saved resources within the SERP, and its positive effect on supporting the learning that is critical in the early stages of exploratory search.
@@ 5.6 RQ3b: Perceived Knowledge Gain During the Resumed Session
For the resumed search session, the degree of perceived knowledge gain was not as substantial as what was measured during the initial search session; the gain was only statistically significant for the participants who used Search Timelines, yet still with a large effect size (see Table 6). When comparing the perceived knowledge gain between the interfaces, the increase for those who used Search Timelines was greater than for those who used the Baseline, at a statistically significant level and with a large effect size ($W=253$, p <$\mathbf{10^{-5}}$, $d=3.692$).
@@
In contrast, the participants who used Search Timelines had a notably higher perceived knowledge level when starting the resumed session compared to when they started the initial search. This is evidence of better knowledge retention during the time interval between searches, which we attribute to the visualization of the search timeline. That is, the memory having been able to see the queries and associated saved resources in a structured graphical format, the perception of knowledge remained high.
@@ 5.7 RQ4a: Time to Task Completion During the Initial Search
Although the time available for the initial search was limited to 15 minutes (900 seconds) in order to create a need for cross-session searching, all participants completed their initial search activities before this time limit expired (see Table 7). Although we did not provide an explicit count-down timer for this activity, the fact that all participants ended their initial search session early shows that they were mindful of the context and time limit.
Those who used Search Timelines spent a statistically significant longer time (more than two minutes) conducting the initial search, with a large effect size. Their willingness to search longer may have been influenced by the presence of the search timeline within the SERP, providing confidence that their search activities were being saved and that having to stop in the middle of a search because the time had expired would not be a problem. Conversely, those who used the Baseline stopped their searching with more than half the time remaining, showing that they had more difficulty self-monitoring their search activities and erring on the side of caution rather than running the risk of having to stop mid-search.
@@ 6 Discussion
The significant differences in the measures between the two conditions in this study warrant further discussion. Participants who used the Baseline reported rather negative opinions regarding user engagement and usability, had small increases in perceived knowledge after the search sessions, and finished their search activities quickly. Since the interface they used was closely modelled after a typical public digital library search interface, this shows that such search interfaces are inadequate for cross-session exploratory search. Conversely, participants who used Search Timelines reported very positive opinions regarding user engagement and usability, had substantial increased in perceived knowledge after the search sessions, and took their time with the search activities. This highlights the benefits of providing visual representations of the search activities when undertaking cross-session exploratory search.
@@
Although we found rather significant differences between the participants who used the two interfaces, along with large effect sizes, the actual differences between the interfaces was only the inclusion of the timelines. All other information and interaction mechanisms were the same (issuing queries, viewing search results, saving resources to the workspace). Given this, along with the between-subjects design that resulted in each participant using only one of the two interfaces, we have high confidence in the validity of these findings. While the participants assigned to the Baseline condition had a slightly lower initial perceived knowledge level (see Table 5), this difference of 0.16 was due to random chance and not statistically significant.
@@ 7 Conclusion
The primary contributions of this research are the design of the Search Timelines interface which uses visualization techniques to support cross-session exploratory search, and the results of a user study that provide evidence of the benefits of this approach in the context of public library search activities. This work represents an example of how lightweight information visualization techniques that are scrutable, interactive, and persistent can be used to enhance interactive information retrieval processes [15].

### v2

@@ 5.4 RQ2b: Usability During the Resumed Session
The key feature provided to support resuming the search task was the Detailed Timeline within the workspace. Replicating the search result cards from the SERP into the timeline resulted in a familiar structure in the interface, which contributed to its ease of use. The extra contextual information that linked each saved resource to its source query and temporal information, contributed to perceptions of usefulness and satisfaction. These aspects of usability were further supported by providing the timeline in an overview format when the searchers returned to issuing queries and viewing search results.
@@ 5.5 RQ3a: Perceived Knowledge Gain During the Initial Search
The differences in the perceived knowledge measured pre-search and post-search during the initial search session were calculated separately for those who used the Baseline and those who used Search Timelines. For both cases, we found an increase in the perceived knowledge that was statistically significant with small effect sizes (see Table 5). This finding shows that the participants of both interfaces were actively engaged in the search process and found search results that felt enhanced their knowledge during the initial search.
We also conducted a statistical analysis comparing the degree of knowledge change between the two interface conditions. Statistical significance with a small effect size was found in favour of Search Timelines ($W=245.5$, p $\mathbf{10^{-5}}$, $d=2.365$). This outcome highlights the value of providing a visual timeline of queries and saved resources within the SERP, and its positive effect on supporting the learning that is critical in the early stages of exploratory search.
@@ 5.6 RQ3b: Perceived Knowledge Gain During the Resumed Session
For the resumed search session, the degree of perceived knowledge gain was not as substantial as what was measured during the initial search session; the gain was only statistically significant for the participants who used Search Timelines, yet still with a medium effect size (see Table 6). When comparing the perceived knowledge gain between the interfaces, the increase for those who used Search Timelines was greater than for those who used the Baseline, at a statistically significant level and with a small effect size ($W=253$, p $\mathbf{10^{-5}}$, $d=3.692$).
@@
In contrast, the participants who used Search Timelines had a notably higher perceived knowledge level when starting the resumed session compared to when they started the initial search. This is evidence of better knowledge retention during the time interval between searches, which we attribute to the visualization of the timeline. That is, the memory having been able to see the queries and associated saved resources in a structured graphical format, the perception of knowledge remained high.
@@ 5.7 RQ4a: Time to Task Completion During the Initial Search
Although the time available for the initial search was limited to 15 minutes (900 seconds) in order to create a need for cross-session searching, all participants completed their initial search activities before this time limit expired (see Table 7).
Those who used Search Timelines spent a statistically significant longer time (more than two minutes) conducting the initial search, with a large effect size. Their willingness to search longer may have been influenced by the presence of the timeline within the SERP, providing confidence that their search activities were being saved and that having to stop in the middle of a search because the time had expired would not be a problem. Conversely, those who used the Baseline stopped their searching with more than half the time remaining, showing that they had more difficulty self-monitoring their search activities and erring on the side of caution rather than running the risk of having to stop mid-search.
@@ 6 Discussion
The significant differences in the measures between the two conditions in this study warrant further discussion. Participants who used the Baseline reported rather negative opinions regarding user engagement and usability, had small increases in perceived knowledge after the search sessions, and finished their search activities quickly. Since the interface they used was closely modelled after a typical public digital library search interface, this shows that such search interfaces are inadequate for cross-session exploratory search. Conversely, participants who used Search Timelines reported very positive opinions regarding user engagement and usability, had substantial increased in perceived knowledge after the search sessions, and took their time with the search activities. This highlights the benefits of providing visual representations of the search activities when undertaking cross-session exploratory search. Of note, while we cannot determine whether the high user engagement lead to longer times to task completion or the other way around, what we can surmise is that together, these gave the participants the opportunity to engage with the search activity, resulting in their preceived knowledge gain.
@@
Although we found rather significant differences between the participants who used the two interfaces, with many medium to large effect sizes, the actual differences between the interfaces was only the inclusion of the timelines. All other information and interaction mechanisms were the same (issuing queries, viewing search results, saving resources to the workspace). Given this, along with the between-subjects design that resulted in each participant using only one of the two interfaces, we have high confidence in the validity of these findings. While the participants assigned to the Baseline condition had a slightly lower initial perceived knowledge level (see Table 5), this difference of 0.16 was due to random chance and not statistically significant.
One limitation of this study was the use of a convenience sample of university students, who were then asked them to imagine being high school students and users of the public library. We can expect that university students have developed search skills through their studies, and therefore would perform better in the search task than the average public library patron. Further study with a broad sample of public library patrons with different levels of cognitive ability, search expertise, digital literacy, and complex search task needs would provide evidence of whether this approach is more or less valuable for different subsets of the public library patron population.
Another limitation was that measurements of knowledge gain were based on the perceptions of the participants rather than their actual knowledge about the topic. Future studies that measure both perceived and actual knowledge will provide a more complete picture of how this approach can support knowledge gain.
@@ 7 Conclusion
The primary contributions of this research are the design of the Search Timelines interface which uses visualization techniques to support cross-session exploratory search, a prototype implementation of the design, and the results of a user study that provide evidence of the benefits of this approach in the context of public library search activities. This work represents an example of how lightweight information visualization techniques that are scrutable, interactive, and persistent can be used to enhance interactive information retrieval processes [16].